# Procountable groups are not classifiable by countable structures

### v3

@@ Abstract.
We prove that topological isomorphism on $2$-nilpotent procountable groups is not classifiable by countable structures, in the sense of descriptive set theory. In fact, the equivalence relation $\ell_{\infty}$ expressing that two sequences of reals have a bounded difference is Borel reducible to it. This marks substantial progress on an open problem of Kechris, Nies and Tent (2018): to determine the exact complexity of the isomorphism relation among all non-archimedean Polish groups.
@@ 1. Introduction
Our work marks considerable progress on an open problem of Kechris et al. KNT pertaining to their programme. They asked at the end of their introduction for the complexity of isomorphism on the class of all non-archimedean Polish groups, observing that this relation is analytic, and GI is Borel reducible to it. By our main result, Theorem 1.2 below, the isomorphism relation on the Borel class of 2-nilpotent procountable groups of exponent a fixed odd prime $p$ already has a complexity much higher than GI.
@@
Procountable groups are of interest because they generalise both profinite and discrete groups, and on the other hand they are precisely the non-archimedean groups that have small invariant neighbourhoods (SIN). (See Pestov:99 for a discussion of this important class of topological groups.) It is easily seen that a group of the form $\prod_{n}G_{n}$, and hence each procountable group, is isomorphic to a closed subgroup of $\mathrm{Sym}(\mathbb{N})$. For background and further equivalent characterisations, see e.g., Malicki2016 . We will verify in 3.3 that the procountable closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a Borel set, using that the procountable groups are the non-archimedean groups with a neighbourhood basis of the neutral element consisting of normal open subgroups (see, e.g., [GX, , Theorem 3.3] or [Malicki2016, , Lemma 2]). The procountable groups satisfying some law (such as being 2-nilpotent) also form a Borel set: from a closed subgroup $G$ of $\mathrm{Sym}(\mathbb{N})$, using [Kechris, , Th. 12.13] one can in a Borel way determine a countable dense subgroup $D$, and it satisfies to require that $D$ satisfies the law.
Since each countable discrete group is procountable, by the aforementioned result of Mekler Mekler1985 , GI is reducible to the isomorphism relation on 2-nilpotent procountable groups of exponent $p$. It follows that the topological isomorphism relation of procountable groups is a complete analytic set and hence non-Borel.
@@
Fix an odd prime $p$. The equivalence relation $\ell_{\infty}$ is Borel reducible to the topological isomorphism relation on 2-nilpotent procountable groups of exponent $p$. Consequently, the topological isomorphism relation on such procountable groups is not Borel reducible to any orbit equivalence relation induced by a Borel action of a Polish group. In particular, it is not Borel reducible to GI.
@@
For an odd prime $p$, let $\mathcal{G}_{2,p}$ be the Borel class of all $2$-nilpotent procountable groups of exponent $p$. There is no assignment $G\mapsto\mathcal{N}_{G}$ such that for each $G\in\mathcal{G}_{2,p}$, $\mathcal{N}_{G}$ is a countable set of open subgroups of $G$ that form a neighborhood basis of $1_{G}$, and this assignment is Borel and isomorphism invariant in the sense of [KNT, , Theorem 3.1].
@@
    Section 4 reduces this uniform homeomorphism relation to the topological isomorphism relation on the Borel space of procountable nilpotent groups of class $2$ and exponent $p$, for $p$ an odd prime.
@@
Some background on the program of KNT . For context we summarize other results known to us that pertain to the program. One class considered in KNT is the class of compact non-archimedean Polish groups, or equivalently, the profinite groups. It was shown that its isomorphism relation is Borel bi-reducible with GI. This implies that this isomorphism relation is a complete analytic set, and hence non-Borel. One can also consider the subclass of abelian profinite groups. Due to Pontryagin duality, the isomorphism relation between abelian profinite groups is Borel bi-reducible with the isomorphism relation between countable torsion abelian groups. The latter relation is classified by Ulm invariants; it is also a complete analytic set and hence non-Borel (see [FriedmanStanley1989, , (2.1)]). In terms of Borel reducibility, it is strictly above the smooth equivalence relation and strictly below GI. In particular, it is not above the equivalence relation $E_{0}$. The isomorphism relation of oligomorphic groups was shown to be essentially countable in NiesSchlichtTent2022 ; it is at present an open question whether this relation is smooth (i.e., real numbers can be assigned as invariants); for partial results on this problem, see NP_oligo ; reconstruction2 ; in particular in reconstruction2 it was shown that under the assumption of weak elimination of imaginaries the relation is smooth. The isomorphism relation of extremely amenable non-archimedean Polish groups was studied in EGL , where it is shown that the setwise equality of infinite sequences of reals is Borel reducible to it. Work in progress of Ruiwen Li, in conjunction with EGL , would show that in fact GI is reducible to it.
@@
Countable groups are up to isomorphism the discrete closed subgroups $G$ of $\mathrm{Sym}(\mathbb{N})$, and being discrete is Borel because it means that the identity permutation is isolated in $G$. If a class of discrete subgroups of $\mathrm{Sym}(\mathbb{N})$ is Borel, its isomorphism relation is Borel reducible to GI. The question is whether conversely, GI reduces to the isomorphism relation on the class. Mekler Mekler1985 (as noted in [FriedmanStanley1989, , 2.3]) established this for the countable groups that are nilpotent of class $2$, and of exponent a fixed odd prime. His method of coding graphs into such groups will be instrumental for our main result, Theorem 1.2. Paolini and Shelah PS showed it for countable torsion-free abelian groups. Most recently, Gao and Li GL did so for countable omnigenous locally finite groups (an uncountable class of groups, each generalising Hall’s universal locally finite group).
@@
We note that our original construction for the second step111see Version 1 of the arxiv preprint, available at arxiv.org/2512.12256.v1 used as the group $G_{n}$ in the inverse system $(G_{n},p_{n})_{n\in\mathbb{N}}$ the free Coxeter group generated by $V_{n}$, namely, the group freely generated by involutions $g_{v}$ for $v\in V_{n}$. That approach does not yield nilpotent-2 groups. However, it is somewhat simpler to verify, because the only elements of order 2 in $G_{n}$ are the conjugates of the generators, so that the elements of $V_{n}$ up to conjugacy can be recovered from $G_{n}$.
@@ 3.2. Proω categories
4. Borel reduction of uniform homeomorphism to isomorphism of 2-nilpotent procountable groups
In this section we implement the second step of our proof strategy toward a proof of our Main Result 1.2, as laid out in the introduction. In 4.7 we will exhibit a Borel reduction from uniform homeomorphism on Polish ultrametric spaces to isomorphism of nilpotent-2 procountable groups of exponent $p$, where $p$ is an odd prime.
4.1. The 2-nilpotent groups $L(X)$
@@
We will only use Mekler’s construction for a very simple infinite graph: the one such that each connected component has size 2. (While we will no use it, we note that this graph is “nice” in the sense of Mekler, cf. [Mekler1985, , Definition 1.1].)
Notation 4.3.
@@
We may assume that $c=\prod_{r\in C}{x_{r}}^{{\alpha}_{r}}$ and $d=\prod_{s\in D}{x_{s}}^{{\alpha}_{s}}$ as in 4.1(ii). If $r\in\mathbb{N}\setminus C$ we let $\alpha_{r}=0$, and if $s\in\mathbb{N}\setminus D$ we let $\beta_{s}=0$. Suppose there is no even $i\in\mathbb{N}$ such that $\langle\underline{c},\underline{d}\rangle=\langle\underline{x}_{i},{\underline{x}}_{i+1}\rangle$. We show that $\underline{c}=\underline{d}^{\lambda}$ for some $\lambda\in\mathbb{N}$.
By assumption there is no $k\in\mathbb{N}$ such that both $C$ and $D$ are contained in $\{2k,2k+1\}$. Note that neither $C$ nor $D$ can be a singleton: if (say) $C=\{r\}$, then there is $s\in D$ such that $r^{*}\neq s^{*}$. Then $\alpha_{r}\beta_{s}\neq 0$ while $\alpha_{s}\beta_{r}=0$, which contradicts the assumption that $c,d]\in Z(L)$ because of 4.2.
@@
$L(X)=N(X)/\langle\{[a_{v},b_{v}]\colon v\in X\}\rangle.$
Given a set $X$, for $v\in X$ we will interpret $a_{v}$ and $b_{v}$ in $L(X)$. Any surjection $q\colon X\to Y$ induces a group epimorphism $\widehat{q}\colon L(X)\to L(Y)$ given by mapping the equivalence class of $a_{v}$ to the equivalence class $a_{q(v)}$, and similarly for $b_{v}$ and $b_{q(v)}$.
An arbitrary bijection $h\colon X\to\mathbb{N}$ induces an isomorphism $L(X)\to L$ (for $L$ as in 4.3) by sending $a_{v}$ to the equivalence class of $x_{2h(v)}$ and $b_{v}$ to the equivalence class of $x_{2h(v)+1}$ given by the presentation in ($\diamond$4.1). So we can apply Lemma 4.4 to $L(X)$ as well. This lemma can now be used to turn certain group epimorphism $L(X)\to L(Y)$ into surjections $X\to Y$ in a functorial manner. This will be needed to eventually verify that the reduction for the second step (4.7 below) works.
@@
Let $\mathcal{L}$ be the category of groups of the form $L(X)$ with morphisms the epimomorphisms $g\colon L(X)\to L(Y)$ such that for each $v\in X$, the rank of $\langle\underline{g(a_{v})},\underline{g(b_{v})}\rangle$ as a subgroup of $L(Y)/Z(L(Y))$ equals $2$. A functor $F$ from $\mathcal{L}$ to the category of countable sets with mappings between them is declared by sending $L(X)$ to $X$, and sending a morphism $g:L(X)\to L(Y)$ to the map $p=F(g)\colon X\to Y$ such that
@@
To show that $F$ is well-defined for a morphism $g$ of $\mathcal{L}$ as above, it suffices to apply 4.4 to $L=L(Y)$, $c=g(a_{v}),d=g(b_{v})$, noting that $[c,d]=e$. Thus in $L(Y)/Z(L(Y))$ we have $\langle\underline{c},\underline{d}\rangle=\langle\underline{a}_{y},\underline{b}_{y}\rangle$ for a unique $y\in Y$, and hence $p(v)=y$ is defined. One verifies easily that $F$ is a functor, and that $F(\widehat{q})=q$. ∎
4.2. The reduction for the second step
Definition 4.7.
@@ 4.2. The reduction for the second step
$V_{n}=L(T_{n})\qquad W_{n}=L(U_{n}).$
So we have inverse systems $(V_{n},\widehat{p}_{n})_{n\in\mathbb{N}}$ and $(W_{n},\widehat{q}_{n})_{n\in\mathbb{N}}$ over the category $\mathrm{GrpEpi}$ determining $G_{T}$ and $G_{U}$, respectively (see 4.5).
$r_{n}\colon T_{\phi(n)}\to U_{n}$ and $s_{n}\colon U_{\psi(n)}\to T_{n}$
so that for each $x\in[T]$ and $y\in[U]$ we have
Let $f_{n}=\widehat{r}_{n}$ and $q_{n}=\widehat{s}_{n}$ as defined in 4.5. The pre-morphisms
Conversely, suppose that $G_{T}$ and $G_{U}$ are isomorphic. By 3.7 there are pre-morphisms $(\overline{f},\phi)\colon(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and $(\overline{g},\psi)\colon(W_{n},\widehat{q}_{n})_{n}\to(V_{n},\widehat{p}_{n})_{n}$ that induce an isomorphism $(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and its inverse, respectively, in $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$. Because the pre-morphisms induce inverses of each other, we have
@@
Using the notation of 4.6, the maps on the left hand sides of the equation are morphisms of the category $\mathcal{L}$. We claim that $g_{n}$ and $f_{n}$ are also morphisms of $\mathcal{L}$ for each $n$. We verify it for $g_{n}$, the case of $f_{n}$ being symmetric. We use the left equation in ($\star$4.2). Write $p$ for $p_{\phi(\psi(n)),n}$. For each $v\in T_{\phi(\psi(n)}$, by definition we have $\widehat{p}(\langle a_{v},b_{v}\rangle)=\langle a_{p(v)},b_{p(v)}\rangle$. Thus the rank of $\langle\underline{a}_{v},\underline{b}_{v}\rangle$ cannot be reduced by applying the homomorphism induced by $f_{\psi(n)}$ on the central quotients. So $f_{\psi(n)}$ is in $\mathcal{L}$ by 4.6, and hence $g_{n}$ is in $\mathcal{L}$ by a similar argument. This shows the claim.
@@
Then $\Phi$ and $\Psi$ are uniformly continuous by 2.5. By ($\star$4.2) and 4.6 we have $p_{\phi(\psi(n)),n}=s_{n}\circ r_{\psi(n)}$ and $q_{\psi(\phi(n)),n}=r_{n}\circ s_{\phi(n)}$ for each $n$. Hence $\Psi\circ\Phi$ is the identity on $[T]$, and $\Phi\circ\Psi$ is the identity on $[U]$. So $T]\cong_{u}[U]$ as required. ∎

### v4

@@ Abstract.
We prove that topological isomorphism on procountable groups is not classifiable by countable structures, in the sense of descriptive set theory. In fact, the equivalence relation $\ell_{\infty}$ expressing that two sequences of reals have a bounded difference is Borel reducible to it. This marks substantial progress on an open problem of Kechris, Nies and Tent (2018): to determine the exact complexity of the isomorphism relation among all non-archimedean Polish groups.
@@ 1. Introduction
Our work marks considerable progress on an open problem of Kechris et al. KNT pertaining to their programme. They asked at the end of their introduction for the complexity of isomorphism on the class of all non-archimedean Polish groups, observing that this relation is analytic, and GI is Borel reducible to it. By our main result, Theorem 1.2 below, the isomorphism relation on the Borel class of procountable groups already has a complexity much higher than GI.
@@
Procountable groups are of interest because they generalise both profinite and discrete groups, and on the other hand they are precisely the non-archimedean groups that have small invariant neighbourhoods (SIN). See Pestov:99 for a discussion of this important class of topological groups. It is easily seen that a group of the form $\prod_{n}G_{n}$, and hence each procountable group, is isomorphic to a closed subgroup of $\mathrm{Sym}(\mathbb{N})$. For background and further equivalent characterisations, see e.g., Malicki2016 . We will verify in 3.3 that the procountable closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a Borel set, using that the procountable groups are the non-archimedean groups with a neighbourhood basis of the neutral element consisting of normal open subgroups (see, e.g., [GX, , Theorem 3.3] or [Malicki2016, , Lemma 2]).
Since each countable discrete group is procountable, by a result of Mekler Mekler1985 , GI is reducible to the isomorphism relation on 2-nilpotent procountable groups of exponent $p$. It follows that the topological isomorphism relation of procountable groups is a complete analytic set and hence non-Borel.
@@
Fix an odd prime $p$. The equivalence relation $\ell_{\infty}$ is Borel reducible to the topological isomorphism relation on procountable groups. Consequently, the topological isomorphism relation on procountable groups is not Borel reducible to any orbit equivalence relation induced by a Borel action of a Polish group. In particular, it is not Borel reducible to GI.
@@
Let $\mathcal{C}$ be the Borel class of all procountable groups. There is no assignment $G\mapsto\mathcal{N}_{G}$ such that for each $G\in\mathcal{C}$, $\mathcal{N}_{G}$ is a countable set of open subgroups of $G$ that form a neighborhood basis of $1_{G}$, and this assignment is Borel and isomorphism invariant in the sense of [KNT, , Theorem 3.1].
@@
    Section 4 reduces this uniform homeomorphism relation to the topological isomorphism relation on the Borel space of procountable groups.
@@
Some background on the program of KNT . For context we summarize other results known to us that pertain to the program. One class considered in KNT is the class of compact non-archimedean Polish groups, or equivalently, the profinite groups. It was shown that its isomorphism relation is Borel bi-reducible with GI. This implies that this isomorphism relation is a complete analytic set, and hence non-Borel. One can also consider the subclass of abelian profinite groups. Due to Pontryagin duality, the isomorphism relation between abelian profinite groups is Borel bi-reducible with the isomorphism relation between countable torsion abelian groups. The latter relation is classified by Ulm invariants; it is also a complete analytic set and hence non-Borel (see [FriedmanStanley1989, , (2.1)]). In terms of Borel reducibility, it is strictly above the smooth equivalence relation and strictly below GI. In particular, it is not above the equivalence relation $E_{0}$. The isomorphism relation of oligomorphic groups was shown to be essentially countable in NiesSchlichtTent2022 ; it is at present an open question whether this relation is smooth (i.e., real numbers can be assigned as invariants); for partial results on this problem, see NP_oligo ; reconstruction2 ; in particular in reconstruction2 it was shown that under the assumption of weak elimination of imaginaries the relation is smooth. The isomorphism relation of extremely amenable non-archimedean Polish groups was studied in EGL , where it is shown that GI is Borel reducible to it.
@@
Countable groups are up to isomorphism the discrete closed subgroups $G$ of $\mathrm{Sym}(\mathbb{N})$, and being discrete is Borel because it means that the identity permutation is isolated in $G$. If a class of discrete subgroups of $\mathrm{Sym}(\mathbb{N})$ is Borel, its isomorphism relation is Borel reducible to GI. The question is whether conversely, GI reduces to the isomorphism relation on the class. Paolini and Shelah PS showed it for countable torsion-free abelian groups. Most recently, Gao and Li GL did so for countable omnigenous locally finite groups (an uncountable class of groups, each generalising Hall’s universal locally finite group).
@@ 3.2. Proω categories
4. Borel reduction of uniform homeomorphism to isomorphism of procountable groups
In this section we implement the second step of our proof strategy toward a proof of our Main Result 1.2, as laid out in the introduction. Definition 4.1 will provide a Borel reduction from uniform homeomorphism on Polish ultrametric spaces to isomorphism of procountable groups.
4.1. The groups $L(V)$
For a countably infinite set $V$, we define the free product of cyclic groups of size $2$ with generators in $V$ by
$L(V)=\langle s\mid s^{2}=e\ \ (s\in V)\rangle.$
(This group is also known as the free Coxeter group of rank $\aleph_{0}$.) By the general theory of free products of group (see, e.g., [LyndonSchupp1977, , IV.4.7]), each involution is conjugate to a generator, and no two generators are conjugate. If $r\colon V\to W$, then we have a map $\widehat{r}\colon L(V)\to L(W)$ induced by $v\mapsto r(v)$.
4.2. The reduction for the second step
Definition 4.1.
[…]
Theorem 4.2.
Proof.
$V_{n}=L(T_{n})\qquad W_{n}=L(U_{n}).$
So we have inverse systems $(V_{n},\widehat{p}_{n})_{n\in\mathbb{N}}$ and $(W_{n},\widehat{q}_{n})_{n\in\mathbb{N}}$ over the category $\mathrm{GrpEpi}$ determining $G_{T}$ and $G_{U}$, respectively (see 4.1).
$r_{n}\colon T_{\phi(n)}\to U_{n}$ and $s_{n}\colon U_{\psi(n)}\to T_{n}$
so that for each $x\in[T]$ and $y\in[U]$ we have
Let $f_{n}=\widehat{r}_{n}$ and $q_{n}=\widehat{s}_{n}$ as defined in Subsection 4.1. The pre-morphisms
Conversely, suppose that $G_{T}$ and $G_{U}$ are isomorphic. By 3.7 there are pre-morphisms $(\overline{f},\phi)\colon(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and $(\overline{g},\psi)\colon(W_{n},\widehat{q}_{n})_{n}\to(V_{n},\widehat{p}_{n})_{n}$ that induce an isomorphism $(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and its inverse, respectively, in $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$. By virtue of being pre-morphisms, the following diagrams are commutative for each $k>n$:
Because the pre-morphisms induce inverses of each other, we have
Let $r_{n}\colon\subseteq T_{\phi(n)}\to U_{n}$ be defined by
$r_{n}(t)=u$ if $f_{n}(t)$ is conjugate in $W_{n}$ to $u$.
(Note here that $f_{n}(t)^{2}=e$, so $f_{n}(t)$ is conjugate to a unique generator, or it equals $e$. The second case could prima facie happen, so $r_{n}$ could be a partial map, as indicated in the notation.) Similarly, let $s_{n}\colon\subseteq U_{\psi(n)}\to T_{n}$ be defined by
$s_{n}(u)=t$ if $g_{n}(u)$ is conjugate in $V_{n}$ to $t$.
We claim that $s_{n}\circ r_{\psi(n)}=p_{\phi(\psi(n)),n}$ for each $n$. To verify this, let $t\in T_{\phi(\psi(n))}$. By ($\star$5.2) we have that $g_{n}(f_{\psi(n)}(t))\neq e$. So
$f_{\psi(n)}(t)\sim_{W_{\psi(n)}}r_{\psi(n)}(t)\in U_{\psi(n)}$
where $\sim_{L}$ denotes conjugacy in a group $L$. So by ($\star$5.2) again
$\widehat{p}_{\phi(\psi(n)),n}(t)=g_{n}(f_{\psi(n)}(t))\sim_{V_{n}}g_{n}(r_{\psi(n)}(t))\sim_{V_{n}}s_{n}(r_{\psi(n)}(t)).$
So $p_{\phi(\psi(n)),n}=s_{n}\circ r_{\psi(n)}$. Similarly, $q_{\psi(\phi(n)),n}=r_{n}\circ s_{\phi(n)}$. Since the compositions are onto and total, each map $r_{n}$ and $s_{n}$ is onto. Thus the maps $r_{n}$ and $s_{n}$ are also total (composing a partial after an onto map yields a partial map).
Corresponding to the diagrams in EpiGrp above, we have commutative diagrams of onto maps between sets
Hence $T]\cong_{u}[U]$ by 2.5, as required.
5. Strengthening to 2-nilpotent groups of exponent $p$
Strengthening the previous result, this section establishes a Borel reduction from uniform homeomorphism on Polish ultrametric spaces to isomorphism of nilpotent-2 procountable groups of exponent $p$, where $p$ is a fixed odd prime. We only need to modify the second step.
5.1. The 2-nilpotent groups $H(X)$
@@
We will only use Mekler’s construction for a very simple infinite graph: the one such that each connected component has size 2.
Notation 5.3.
@@
We may assume that $c=\prod_{r\in C}{x_{r}}^{{\alpha}_{r}}$ and $d=\prod_{s\in D}{x_{s}}^{{\alpha}_{s}}$ as in 5.1(ii). If $r\in\mathbb{N}\setminus C$ we let $\alpha_{r}=0$, and if $s\in\mathbb{N}\setminus D$ we let $\beta_{s}=0$. Suppose there is no even $i\in\mathbb{N}$ such that $\langle\underline{c},\underline{d}\rangle=\langle\underline{x}_{i},{\underline{x}}_{i+1}\rangle$. We show that $\underline{c}=\underline{d}^{\lambda}$ for some $\lambda\in\mathbb{N}$.
By assumption there is no $k\in\mathbb{N}$ such that both $C$ and $D$ are contained in $\{2k,2k+1\}$. Note that neither $C$ nor $D$ can be a singleton: if (say) $C=\{r\}$, then there is $s\in D$ such that $r^{*}\neq s^{*}$. Then $\alpha_{r}\beta_{s}\neq 0$ while $\alpha_{s}\beta_{r}=0$, which contradicts the assumption that $c,d]\in Z(L)$ because of 5.2.
@@
$H(X)=N(X)/\langle\{[a_{v},b_{v}]\colon v\in X\}\rangle.$
Given a set $X$, for $v\in X$ we will interpret $a_{v}$ and $b_{v}$ in $H(X)$. Any surjection $q\colon X\to Y$ induces a group epimorphism $\widehat{q}\colon H(X)\to H(Y)$ given by mapping the equivalence class of $a_{v}$ to the equivalence class $a_{q(v)}$, and similarly for $b_{v}$ and $b_{q(v)}$.
An arbitrary bijection $h\colon X\to\mathbb{N}$ induces an isomorphism $H(X)\to L$ (for $L$ as in 5.3) by sending $a_{v}$ to the equivalence class of $x_{2h(v)}$ and $b_{v}$ to the equivalence class of $x_{2h(v)+1}$ given by the presentation in ($\diamond$5.1). So we can apply Lemma 5.4 to $H(X)$ as well. This lemma can now be used to turn certain group epimorphism $H(X)\to H(Y)$ into surjections $X\to Y$ in a functorial manner. This will be needed to eventually verify that the reduction for the second step (4.1 below) works.
@@
Let $\mathcal{L}$ be the category of groups of the form $H(X)$ with morphisms the epimomorphisms $g\colon H(X)\to H(Y)$ such that for each $v\in X$, the rank of $\langle\underline{g(a_{v})},\underline{g(b_{v})}\rangle$ as a subgroup of $H(Y)/Z(H(Y))$ equals $2$. A functor $F$ from $\mathcal{L}$ to the category of countable sets with mappings between them is declared by sending $H(X)$ to $X$, and sending a morphism $g:H(X)\to H(Y)$ to the map $p=F(g)\colon X\to Y$ such that
@@
To show that $F$ is well-defined for a morphism $g$ of $\mathcal{L}$ as above, it suffices to apply 5.4 to $L=H(Y)$, $c=g(a_{v}),d=g(b_{v})$, noting that $[c,d]=e$. Thus in $H(Y)/Z(H(Y))$ we have $\langle\underline{c},\underline{d}\rangle=\langle\underline{a}_{y},\underline{b}_{y}\rangle$ for a unique $y\in Y$, and hence $p(v)=y$ is defined. One verifies easily that $F$ is a functor, and that $F(\widehat{q})=q$. ∎
5.2. The new reduction for the second step
Given a pruned tree $T$, let the sets $T_{n}$ be as before, and let
@@ 5.2. The new reduction for the second step
Let $T$ and $U$ be pruned trees on $\mathbb{N}$ such that $T_{1}$ and $U_{1}$ are infinite. Then $T]\cong_{u}[U]$ if and only if $H_{T}$ is topologically isomorphic to $H_{U}$.
@@
As before, let the maps $p_{n}\colon T_{n+1}\to T_{n}$ and $q_{n}\colon U_{n+1}\to U_{n}$ be given by the predecessor maps on the trees $T$ and $U$, respectively. Now write
$V_{n}=H(T_{n})\qquad W_{n}=H(U_{n}).$
So we have inverse systems $(V_{n},\widehat{p}_{n})_{n\in\mathbb{N}}$ and $(W_{n},\widehat{q}_{n})_{n\in\mathbb{N}}$ over the category $\mathrm{GrpEpi}$ determining $H_{T}$ and $H_{U}$, respectively (see 5.5).
If $T]\cong_{u}[U]$ via a uniform homeomorphism $\Phi$, then $H_{T}\cong H_{U}$ as before. Now suppose that $H_{T}$ and $H_{U}$ are isomorphic. By 3.7 there are pre-morphisms $(\overline{f},\phi)\colon(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and $(\overline{g},\psi)\colon(W_{n},\widehat{q}_{n})_{n}\to(V_{n},\widehat{p}_{n})_{n}$ that induce an isomorphism $(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and its inverse, respectively, in $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$. Because the pre-morphisms induce inverses of each other, we have
@@
Using the notation of 5.6, the maps on the left hand sides of the equation are morphisms of the category $\mathcal{L}$. We claim that $g_{n}$ and $f_{n}$ are also morphisms of $\mathcal{L}$ for each $n$. We verify it for $g_{n}$, the case of $f_{n}$ being symmetric. We use the left equation in ($\star$5.2). Write $p$ for $p_{\phi(\psi(n)),n}$. For each $v\in T_{\phi(\psi(n)}$, by definition we have $\widehat{p}(\langle a_{v},b_{v}\rangle)=\langle a_{p(v)},b_{p(v)}\rangle$. Thus the rank of $\langle\underline{a}_{v},\underline{b}_{v}\rangle$ cannot be reduced by applying the homomorphism induced by $f_{\psi(n)}$ on the central quotients. So $f_{\psi(n)}$ is in $\mathcal{L}$ by 5.6, and hence $g_{n}$ is in $\mathcal{L}$ by a similar argument. This shows the claim.
@@
Then $\Phi$ and $\Psi$ are uniformly continuous by 2.5. By ($\star$5.2) and 5.6 we have $p_{\phi(\psi(n)),n}=s_{n}\circ r_{\psi(n)}$ and $q_{\psi(\phi(n)),n}=r_{n}\circ s_{\phi(n)}$ for each $n$. Hence $\Psi\circ\Phi$ is the identity on $[T]$, and $\Phi\circ\Psi$ is the identity on $[U]$. So $T]\cong_{u}[U]$ as required. ∎